OpenStudy (anonymous):

Which of the following is a polynomial with roots 5, 7, and −8? f(x) = x3 − 35x2 − 56x + 61 f(x) = x3 − 35x2 − 61x + 280 f(x) = x3 − 4x2 − 35x + 56 f(x) = x3 − 4x2 − 61x + 280

OpenStudy (anonymous):

@MrNood @OregonDuck

oregonduck (oregonduck):

f(x) = (x-5)(x-7)(x+8) = (x-5)(x^2+x-56) = x^3+x^2-56x-5x^2-5x+280 f(x) = x^3-4x^2-61x+280 So the answer is d

OpenStudy (anonymous):

how did you do that so fast

OpenStudy (anonymous):

are you super man?

oregonduck (oregonduck):

cause this stuff is easy

oregonduck (oregonduck):

i can do pre calculus in 7th grade

oregonduck (oregonduck):

i am only 12 fyi

OpenStudy (anonymous):

so you are super man

OpenStudy (anonymous):

mind helping me with one more

OpenStudy (mrnood):

however - @OregonDuck all oyu have given the poster is the knowledge that YOU can solve this question what iwould be truly helpful would be to show HOW to solve it - then let HIM do the work you learn better by DOING than by watching

oregonduck (oregonduck):

sure :)

OpenStudy (anonymous):

Which of the following is a polynomial with roots 4, 2i, and −2i? f(x) = x3 − 4x2 + 8x − 4 f(x) = x3 − 8x2 + 16x − 4 f(x) = x3 − 4x2 + 4x − 16 f(x) = x3 − 4x2 + 8x − 16

OpenStudy (mrnood):

wait!

oregonduck (oregonduck):

WHAT?!?

OpenStudy (mrnood):

if the roots are a,b & c then the factors are (x-a)(x-b)(x-c) if you put your three roots into that formula then mutiply out the 23 brackets you will come to the answer YOURSELF

oregonduck (oregonduck):

Work backwards for each root to get the factor.. (x - 4)(x - 2i)(x + 2i) Distribute.. (x - 4)(x² - 4i²) = (x - 4)(x² + 4) [Noting that i² = -1] = x³ - 4x² + 4x - 16 [A]

OpenStudy (mrnood):

if the roots are a,b & c then the factors are (x-a)(x-b)(x-c) if you put your three roots into that formula then mutiply out the 3 brackets you will come to the answer YOURSELF

OpenStudy (anonymous):

ya i got that to

oregonduck (oregonduck):

it is a

oregonduck (oregonduck):

wait no it is c sorry i was reading mine post i put a for some reason lol

OpenStudy (mrnood):

@OregonDuck Once again - juist showing that YOU know how to do this type of questin. LET the poster do some of the calculation - he will actually LEARN from that instead of copying

OpenStudy (anonymous):

im working it out as well mrnood

oregonduck (oregonduck):

he will learn my failing the dba or segment exams cause you need your notes and memory to help you

OpenStudy (anonymous):

i am taking these problems down as notes

OpenStudy (anonymous):

so i can learn to do the harder ones

oregonduck (oregonduck):

EXACTLY SEE ^^^^^^^^

OpenStudy (mrnood):

but until you do one yourself you will not be sure that oyu UNDERSTAND how to do it you do not need to learn 'harder' questions - you need to learnt the METHOD - then oyu can do ALL questions

OpenStudy (anonymous):

i did 8 out of the 10 of my practice questions @MrNood but just needed help for these two

oregonduck (oregonduck):

FINE ASK HIM FOR HELP NOT ME AL LEAST I SHOWED MY WORK!!!!!!!

OpenStudy (anonymous):

Please dont fight @OregonDuck and @MrNood

OpenStudy (mrnood):

Don't get in a sulk Just try to help posters to learn rather than going right through to tthe answer

OpenStudy (anonymous):

thank you @MrNood and @OregonDuck

oregonduck (oregonduck):

yeah your welcome

OpenStudy (mrnood):

YOU'RE welcome